To build an interval estimate for the proportion (percentage) of American households who live paycheck-to-paycheck, a random sample 750 full-time workers showed 615 live paycheck-to-paycheck.
15 The point estimate of the population proportion is ________.
A 0.84
B 0.82
C 0.80
D 0.78
16 The standard error of the sample proportion is ________.
A 0.0140
B 0.0175
C 0.0208
D 0.0231
Answer #1

Solution :

Given that,

n = 750

x = 615

15.

Point estimate = sample proportion = = x / n = 615/750=0.82

correct option b

16.

standard error= ((( * (1 - )) / n)

= (((0.82*(1-.82)) /750 )

= 0.0140